a) Buffer solutions with a pH values of around 10 are prepared using sodium carbonate (Na2CO3) and sodium hydrogen carbonate (NaHCO3). What is the pH of a solution of 10.0g each of the two salts in enough water to make 0.250 L of solution? pKa2= 10.33

b) By how much does the pH change when 3.5mL of 6.0 M HCl is added?

c) By how much does the pH of the solution change when 0.92 g of NaOH is added?
Answer #1

Answer:-

(a)-

Given:-

wt. of sodium carbonate (Na2CO3) = 10.0 g

wt. of sodium hydrogen carbonate (NaHCO3) = 10.0 g

volume of buffer solution (V) =  0.250 L

pKa2 value = 10.33

As we know that

molar mass of sodium carbonate (Na2CO3) = 2 molar mass of Na + molar mass of C + 3 molar mass of O

molar mass of sodium carbonate (Na2CO3) = 2 23 + 12 + 3 16

molar mass of sodium carbonate (Na2CO3) = 46 + 12 + 48

molar mass of sodium carbonate (Na2CO3) = 106 g / mol

similarly

molar mass of sodium hydrogen carbonate (NaHCO3) = molar mass of Na + molar mass of H + molar mass of C + 3 molar mass of O

molar mass of sodium hydrogen carbonate (NaHCO3) = 23 + 1 + 12 + 3 16

molar mass of sodium hydrogen carbonate (NaHCO3) = 23 + 1 + 12 + 48

molar mass of sodium hydrogen carbonate (NaHCO3) = 84 g / mol

therefore

No. of moles of sodium carbonate (Na2CO3) = wt. of sodium carbonate (Na2CO3) / molar mass of sodium carbonate (Na2CO3)

No. of moles of sodium carbonate (Na2CO3) = 10.0 g / 106 g / mol

No. of moles of sodium carbonate (Na2CO3) = 0.094 mol

similarly

No. of moles of sodium hydrogen carbonate (NaHCO3) = wt. of ssodium hydrogen carbonate (NaHCO3) / molar mass of sodium hydrogen carbonate (NaHCO3)

No. of moles of sodium hydrogen carbonate (NaHCO3) = 10.0 g / 84 g / mol

No. of moles of sodium hydrogen carbonate (NaHCO3) = 0.119 mol

since we know that

molar concentration of compound = No. of moles of compound (n) / volume of solution (V)

molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = No. of moles of sodium carbonate (Na2CO3) / volume of buffer solution (V)

molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = 0.094 mol / 0.250 L

molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = 0.376 mol /  L = 0.376 M

similarly

molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] = No. of moles of sodium hydrogen carbonate (NaHCO3) / volume of buffer solution (V)

molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] = 0.119 mol / 0.250 L

molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] = 0.476 mol /L = 0.476 M

So according to the Henderson-Hasselbalch equation

pH = pKa + log [salt] /[acid]

since sodium hydrogen carbonate (NaHCO3) = acidic salt i.e acid

sodium carbonate (Na2CO3) = salt

therefore

pH = pKa2 + log [Na2CO3] /[NaHCO3]

pH =10.33 + log (0.376) / (0.476)

pH =10.33 + log (0.376 / 0.476)

pH =10.33 + log (0.7899)

pH =10.33 + ( - 0.1024)

pH =10.33 - 0.1024

pH =10.23 (i.e the answer)

(b)-

Given:-

volume of HCl = 3.5 mL = 3.5 / 1000 = 3.5 10-3 L

molarity of HCl = 6.0 M = 6.0 mol / L

As we know that

molarity of compound = no. of moles of compound (n) / volume of solution (V) in liter (L)

molarity of HCl = no. of moles of HCl (n) / volume of solution (V) in liter (L)

no. of moles of HCl (n) = molarity of HCl volume of solution (V) in liter (L)

no. of moles of HCl (n) = 6.0 mol / L  ​​​​​​​ 3.5 10-3 L

no. of moles of HCl (n) = 21 10-3 mol

no. of moles of HCl (n) = 0.021 mol

Since we know that

No. of moles of sodium carbonate (Na2CO3) = 0.094 mol

No. of moles of sodium hydrogen carbonate (NaHCO3) = 0.119 mol

volume of buffer solution (V) =  0.250 L

when addition of 3.5 mL of 6.0 M HCl i.e 0.021 mol of HCl takes place buffer solution of sodium hydrogen carbonate (NaHCO3) and sodium carbonate (Na2CO3) then no. of moles of sodium carbonate (Na2CO3) is decreased due to reaction between sodium carbonate (Na2CO3) and HCl and no. of moles of sodium hydrogen carbonate (NaHCO3) is increased as follows:-

Na2CO3 + HCl NaHCO3 + NaCl

Initial 0.094 mol 0    0.119 mol 0

Change - 0.021 mol - 0.021 mol + 0.021 mol

therefore

decreased no. of moles of sodium carbonate (Na2CO3) = 0.094 mol - 0.021 mol = 0.073 mol

increased  no. of moles of sodium hydrogen carbonate (NaHCO3) = 0.119 mol + 0.021 mol = 0.140 mol

So

New molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = No. of moles of sodium carbonate (Na2CO3) / volume of buffer solution (V)

New molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = 0.073 mol / 0.250 L

New molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = 0.292 mol /  L = 0.292 M

similarly

New molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] =0.140 mol /0.250 L

New molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] = 0.560 mol /  L = 0.560 M

So according to the Henderson-Hasselbalch equation

pH = pKa + log [salt] /[acid]

pH = pKa2 + log [Na2CO3] /[NaHCO3]

pH =10.33 + log (0.292) / (0.560)

pH =10.33 + log (0.292 / 0.560)

pH =10.33 + log (0.5214)

pH =10.33 + ( - 0.2828)

pH =10.33 - 0.2828

pH =10.05

Change in pH of buffer solution = pH of buffer solution after addition of HCl - Initial pH of buffer solution

Change in pH of buffer solution = 10.05 - 10.23

Change in pH of buffer solution = - 0.18 pH decreased (i.e the answer)

(c)-

Given:-

wt. of NaOH  = 0.92 g

As we know that

molar mass of NaOH =   molar mass of Na + molar mass of O + molar mass of H

molar mass of NaOH =   23 + 16 + 1

molar mass of NaOH =   40 g / mol

Also we know that

No. of moles of NaOH = wt. of NaOH  / molar mass of NaOH

No. of moles of NaOH = 0.92 g  / 40 g / mol

No. of moles of NaOH = 0.023 mol

Since we know that

No. of moles of sodium carbonate (Na2CO3) = 0.094 mol

No. of moles of sodium hydrogen carbonate (NaHCO3) = 0.119 mol

volume of buffer solution (V) =  0.250 L

when addition of 0.92 g of NaOH i.e 0.023 mol of NaOH takes place buffer solution of sodium hydrogen carbonate (NaHCO3) and sodium carbonate (Na2CO3) then no. of moles of sodium hydrogen carbonate (NaHCO3) is decreased due to reaction between sodium hydrogen carbonate (NaHCO3) and NaOH and no. of moles of sodium carbonate (Na2CO3) is increased as follows:-

NaHCO3 + NaOH Na2CO3 + H2O

Initial 0.119 mol 0    0.094 mol 0

Change - 0.023 mol - 0.023 mol + 0.023 mol

therefore

increased no. of moles of sodium carbonate (Na2CO3) = 0.094 mol + 0.023 mol = 0.117 mol

decreased  no. of moles of sodium hydrogen carbonate (NaHCO3) = 0.119 mol - 0.023 mol= 0.096 mol

So

New molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = No. of moles of sodium carbonate (Na2CO3) / volume of buffer solution (V)

New molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = 0.117 mol / 0.250 L

New molar concentration of sodium carbonate (Na2CO3) i.e [Na2CO3] = 0.468 mol /  L = 0.468 M

similarly

New molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] =0.096 mol /0.250 L

New molar concentration of sodium hydrogen carbonate (NaHCO3) i.e [NaHCO3] = 0.384 mol /  L = 0.384 M

So according to the Henderson-Hasselbalch equation

pH = pKa + log [salt] /[acid]

pH = pKa2 + log [Na2CO3] /[NaHCO3]

pH =10.33 + log (0.468) / (0.384)

pH =10.33 + log (0.468 / 0.384)

pH =10.33 + log (1.21875)

pH =10.33 + ( 0.08591)

pH =10.33 + 0.08591

pH =10.42

Change in pH of buffer solution = pH of buffer solution after addition of NaOH - Initial pH of buffer solution

Change in pH of buffer solution = 10.42 - 10.23

Change in pH of buffer solution = 0.19 pH increased (i.e the answer)